# Entanglement Dynamics in Monitored Systems and the Role of Quantum Jumps

### v2

@@ VII.1 Monitored SSH
Figure 10: Steady-State Entanglement Balance for the Monitored Ising chain: Average entanglement gain $\overline{\delta S_{\rm nH}}$ and loss $-\overline{\delta S_{\rm qj}}/\overline{\tau}$, as a function of the average entanglement entropy $S$ and different subsystem sizes $\ell$ (error bars are smaller than the size of the dots). In the weak monitoring phase, Panel (a), both contributions scale logarithmically with $\ell$ (see the inset), suggesting the QJs weakly renormalize the non-Hermitian dynamics. Both terms are independent of subsystem size at strong monitoring (panel b), leading to an area law. Increasing the transverse field (panel c for $h=1$), the gain contribution due to the non-Hermitian dynamics acquires a non-trivial $\ell$ dependence (again scaling logarithmically, see inset). Together with the QJ loss term, this leads to a sub-volume steady-state entanglement absent in the no-click limit. For even larger $h$ (panel d, $h=4$), the non-Hermitian gain term becomes $\ell$ independent, as in the no-click limit. At the same time, the QJ retains a non-trivial scaling (inset), leading to the observed logarithmic scaling.
This might come as a surprise at first since the bare (no-click) non-Hermitian dynamics would show a clear volume-law scaling for this parameter value. However, unlike the no-click limit, our QJ dynamics reach a steady state when gain and loss are balanced, a condition that is reached for much smaller entanglement entropy. QJs, in other words, confine the non-Hermitian dynamics to small values of entanglement entropy, where it is essentially independent of $\ell$. We can, therefore, interpret this result by saying that the effect of QJs here is to renormalize the non-Hermitian dynamics into an effective area-law phase. In Appendix C, we discuss the dynamics from a highly entangled state to show how, in that case, one could retrieve a subsystem size dependence in the non-Hermitian contribution, but only at large values of entropy, far above what the system can explore under the effects of QJs.
From this result, we conclude that for the monitored SSH, the logarithmic growth of the entanglement found in Sec. IV arises mainly due to QJs. The picture is that the main effect of increasing the subsystem size is to decrease the slope of $\Delta S_{\rm qj}$, i.e. decrease the rate at which QJs reduce entanglement thus allowing the system to reach a higher value of entanglement entropy. From this point of view, the growth of $S$ should be understood more as a suppression of entanglement loss, due to the jumps being less effective in disentangling a large subsystem. For large monitoring (Fig 9 (b)), on the other hand, we see that both contributions are almost independent on $\ell$, and their crossing occurs in a regime of small entanglement entropy, leading to an area-law scaling.
VII.2 Monitored Ising Chain
Quite interestingly, the situation is rather different for the Ising chain, as shown in Fig. 10. Let us first consider the weak-monitoring and weak-field regime $\gamma=0.2,\quad h=0.4$ (panel a), corresponding to the log-phase (see Sec. IV).
As we can see, the entanglement gain $\overline{\delta S_{\rm nH}}$ and loss $-\overline{\Delta S_{\rm qj}}/\overline{\tau}$ here show both a non-trivial $\ell$ dependence, which appears compatible with a logarithmic scaling (see inset). We note that the scaling collapse of our data is expected to occur around the crossing point, indicating the steady-state value $S_{\infty}$. In particular, while the entanglement loss due to QJs behaves similarly to the SSH case, decreasing with $\ell$ in absolute value, the entanglement gain due to the renormalized non-Hermitian dynamics, shows a logarithmic scaling with $\ell$, as in the no-click limit at this value of parameters. Based on these results, we conclude that in the weak monitoring phase, the non-Hermitian dynamics are only weakly renormalized by QJs. Suppose we increase the monitoring strength $\gamma=3.0$ (panel b). In that case, the system enters the area law, and we can see that both gain and loss contributions become essentially independent from subsystem size, crossing in a low-entanglement region as expected from an area law.
@@ VII.2 Monitored Ising Chain
Let us now consider the region of weak monitoring and large transverse field, where major deviations from the no-click limit were reported (see discussion in Sec. IV). In particular, we fix $\gamma=0.2$ and scan the transverse field for increasing values of $h>1$, where the no-click evolution predicts an area-law for the entanglement entropy. At the same time, the QJ dynamics still showed a logarithmic scaling. In Fig. 10 (c-d), we repeat our analysis of average entanglement gain and loss for different subsystem sizes $\ell$. For $h=1.0$, we again observe strong $\ell$ dependence in the QJ contribution and the non-Hermitian one. Our steady-state condition (25) predicts a logarithmic scaling for the entanglement entropy, in perfect agreement with the QJ simulation. By breaking down the entanglement content into gain and loss, we can clarify that this log-phase is due to a combined effect of quantum jumps and renormalised non-Hermitian evolution.
@@
Finally, upon increasing the transverse field to $h=4$, we see that the entanglement gain due to the non-Hermitian dynamics becomes essentially $\ell$ independent, in agreement with the no-click limit. On the other hand, the loss due to QJs scales logarithmically with $\ell$ (see inset), which results in a steady-state entanglement showing a log-scaling. This logarithmic phase, as for the weak monitoring of the SSH model, is solely due to the QJs. Our analysis therefore highlights two mechanisms at play behind the deviations from the no-click limit observed in the monitored Ising chain phase diagram for $h>1$. At moderate fields there is a non-trivial renormalization of the non-Hermitian dynamics due to QJs, that leads to a log-scaling of its average gain contribution, as in the weak-monitoring regime of the model. At larger fields instead this renormalisation is washed away, the non-Hermitian dynamics behaves as in the no-click limit and the logarithmic scaling of the entropy arises only from the jumps. While our numerical results does not allow to conclude about the large system size limit, the above analysis suggests that, in absence of a non-trivial non-Hermitian dynamics, the large field log-phase might eventually saturate at large system sizes into an area law, similarly to the weak monitoring phase of the SSH model.
VII.3 Discussion and Implications for Free Fermions MIPT
Our findings for the average entanglement gain and loss have highlighted the concept of renormalized non-Hermitian dynamics. Due to quantum jumps, which effectively reshuffle the initial state onto which the non-Hermitian Hamiltonian act, the effective non-Hermitian dynamics is renormalized with respect to the bare no-click evolution. We have shown that this renormalization is particularly strong in the weak monitoring phase of the SSH chain, where the jumps wash the no-click volume law entirely away. On the other hand, our results for the Ising chain at weak monitoring points towards a much weaker renormalization. Our results, therefore, highlight a fundamental difference between the weak monitoring phase of the Ising and the SSH chain. While both phases feature a logarithmic scaling of the entanglement entropy, we have demonstrated that the origin of this scaling is very different in the two cases. For the SSH, it comes essentially from the quantum jumps, while for the Ising chain, it is a combined effect of jumps and non-Hermitian evolution. The same renormalization effect we have shown to be at play in the monitored Ising chain at intermediate and large field values and can, therefore, naturally explain the departure from the no-click limit.
We conclude with a discussion of the implications of our results for the MIPT of free fermions under quantum jumps. The irrelevance of QJs at weak monitoring in the Ising chain points towards an important role of the non-Hermitian Hamiltonian in this regime. This suggests that the logarithmic phase of entanglement entropy under QJs could be possibly stable in the large system size, as in the no-click case Turkeshi and Schiró (2023). This is consistent with the replica field theory prediction for noisy Majorana fermions under QSD Fava et al. (2023), although we emphasize the differences in measurement protocol and unitary dynamics between the two cases. In particular based on our numerical results we cannot establish whether also in the Ising chain the entanglement scaling is $\ln^{2}(\ell)$. On the other hand our results for the SSH case points toward a strong effect of QJs at weak monitoring. As we have shown the logarithmic phase found numerically is mainly due to QJs. Their contribution is likely to saturate at large system sizes through a mechanism similar to the weak-localization correction Poboiko et al. (2023a).
VIII Conclusions
In this work, we studied the role of quantum jumps on the entanglement dynamics of two models of monitored free fermions, the SSH and Ising chains. In both cases, the no-click limit, corresponding to a deterministic non-Hermitian evolution, was studied in detail and revealed a rich pattern of entanglement transitions. The inclusion of QJ has a somewhat different effect in the two models. On one hand, in the SSH case, the inclusion of jumps qualitatively changes the physics of the no-click limit, washing away the volume-to-area law transition and leaving a sub-volume phase with logarithmic scaling of entanglement entropy, at least for the system sizes considered in this work. On the other hand, for the Ising chain, our results confirm that the effect of stochastic jumps is mainly quantitative, except in specific areas of the phase diagram where deviations from the no-click limit appear.
To better understand the origin of such a different impact of QJs, we have first computed their statistics, particularly their waiting-time distribution. Surprisingly, we have found that at long times, both models acquire a universal Poisson distribution of waiting times, with the same average waiting time $\bar{\tau}$. Then, we introduced an unexplored characterization based on the entanglement gain and loss statistics after or between quantum jumps. The resulting histogram display, particularly for the quantum jump case, a very broad distribution and a typical value pinned close to zero, suggesting that most QJs are not significantly affecting the entanglement controlled by rare jump events.
Using the full statistics of entanglement gain and loss, we have built a stochastic random walk model with partial resetting, which can reproduce the full QJ dynamics of entanglement entropy for both models. This model offers a new light for interpreting the QJ numerics. Indeed, it suggests a natural steady-state condition for the entanglement entropy, given by the balance between gain due to the non-Hermitian evolution and losses due to quantum jumps. Remarkably, this condition accurately reproduced the scaling of entanglement entropy with subsystem size obtained by the full QJ dynamics. Furthermore, it clarified the origin of the different entanglement scaling and the mutual role of jumps and non-Hermitian evolution.
The outcome of this analysis reveals a compelling difference between the two models. Indeed, while for the SSH, the sub-volume phase at weak-monitoring is mainly due to the jumps, which allows larger and larger values of entanglement entropy to be accessed, leading to an overall growth with $\ell$, for the Ising chain, the weak-monitoring phase arises from the interplay of jumps becoming less effective in destroying entanglement and non-Hermitian evolution creating more and more entanglement. Furthermore, our analysis allows us to identify the source of the unexpected logarithmic scaling of the entanglement in the monitored Ising chain at large fields due to the renormalized non-Hermitian evolution dressed by quantum jumps.
An interesting future direction could be to use the statistics of entanglement gain and loss introduced here to study non-integrable, interacting monitored quantum systems under quantum jumps Fuji and Ashida (2020). Here one can expect the volume law phase to be stable, at least up to a critical monitoring rate. One could therefore expect a different statistics of entanglement loss in this case, reflecting the differences in how gaussian and genuine ergodic quantum states respond to quantum measurements.
Acknowledgements.
We thank A. Biella, M. Buchhold, J. Dalibard, M. Dalmonte, R. Fazio, A. Paviglianiti, L. Piroli, A. Romito, P. Sierant, and A. Silva for discussions and collaborations on related topics. We acknowledge computational resources on the Collége de France IPH cluster. X.T. acknowledge DFG under Germany’s Excellence Strategy – Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 – 390534769, and DFG Collaborative Research Center (CRC) 183 Project No. 277101999 - project B01. Note added. During the completion of this work, we become aware of a related manuscript by C. Leung, et al. appearing in the same arXiv posting, studying the stability of the non-Hermitian Hamiltonian for quantum state diffusion C. Leung, et al., To appear .
Appendix A Statistics of the monitored observable
A different perspective on the role of quantum jumps in the monitored dynamics of our two models can be obtained by looking at the statistics of the monitored observable, in this case, the histogram of the local density. As discussed in Sec. III of the main text, this provides information that is not contained in the average dynamics described by the Lindbladian, which only captures the average value (first moment) of the histogram.
To appreciate this point in Fig. 11 we plot for the monitored SSH the statistics of the particle density in the $A$ sublattice for increasing values of the monitoring rate $\gamma$. Given the choice of our initial state, this histogram at time $t=0$ is peaked at $n_{A}=0$, but at long times we see a redistribution due to the interplay of non-Hermitian dynamics and jumps. For weak monitoring, we see that the long-time limit is an almost flat distribution where all measurement outcomes are almost equally probable, except for the region near $n_{A}=1$, which features a small peak reminiscent of the fact that jumps in the SSH model gets rid of state with empty sites in the $A$ sublattice. For comparison, we show the average density predicted by the non-Hermitian evolution as a dashed line.
Figure 11: Statistics of the particle density for the SSH (left panels), corresponding to the sublattice $A$, and for the Ising chain (right panels). Upon increasing the monitoring rate we see the emergence of a strong bimodality, with one of the two peaks that appear in correspondence of the non-Hermitian evolution result.
As the monitoring strength increases, we see that the most likely outcome is always $n_{A}=1$. However, there is a secondary peak that appears near $n_{A}=0$. For strong monitoring, when the system is in the area of law, this secondary peak is very close to the value predicted by the non-Hermitian Hamiltonian. This analysis confirms, therefore, that, at large monitoring, the influence of the non-Hermitian evolution is present in the full QJ dynamics. It also explains that the entanglement entropy shows an area law scaling in both cases since the long-time steady-state limit at large $\gamma$ are two CDW states that only have $O(1)$ entanglement.
In the right panels of Fig. 11, we plot the particle density (or local transverse magnetization) statistics for the Ising chain. Here, we see a different behavior at weak monitoring; a gap appears in the histogram of particle density in correspondence with the values $n=0$, which is gradually filled with increasing $\gamma$. As the monitoring rate is further increased, a bimodal structure emerges, as already observed in Ref. Biella and Schiró (2021), in correspondence with the value expected from the non-Hermitian dynamics (dashed line).
We emphasize that to appreciate this point, we need to look at the full histogram of particle density since the average value described by the Lindblad master equation only illustrates that the system wants to reach a uniform steady state.
[…]
Appendix B System size scaling of entanglement loss
In the main text we introduced a classical stochastic model for the entanglement entropy dynamics, which builds upon the entanglement gain and loss statistics. In particular, the role of QJs is to induce an entanglement loss that, in the stationary state, is balanced by the gain provided by the non-Hermitian evolution, as described in Eq. (24) of the main text. This equation divides the entanglement loss by the average waiting time $\overline{\tau}$, which considers the instantaneous nature of QJs. As we have seen in Sec. V, the average waiting time scales as $1/L$, which raises the question of the stability of the steady-state condition in the large system size limit. In this Appendix, we provide evidence supporting the statement that the average entanglement loss due to the jump scales is $1/L$, so the ratio with the averaging waiting time remains finite when $L\rightarrow\infty$. In Fig. 12, we plot, for the monitored Ising chain at a representative value of the parameters, the average entanglement loss divided by the waiting time, $\overline{\Delta S_{\rm qj}}(S,\ell)/\overline{\tau}$ as a function of $S,\ell$ and different system sizes. We see that the data corresponding to different sizes collapsing onto each other for each subsystem sizes $\ell$, which demonstrate this invariance of the ratio $\overline{\Delta S_{\mathrm{qj}}}/\overline{\tau}$.
Figure 13: Analysis of trajectories obtained from a highly entangled initial state in the SSH model. Panel (a) : Entanglement Dynamics in the monitored SSH chain under QJ dynamics with an initial state highly entangled (namely a steady state of the no-click limit). Panel (b) : Conditioned version of the distribution $\Delta S_{\rm qj}$ with respect to the entanglement when the jumps is happening. Panel (c) : First moment of the conditioned distribution $\Delta S_{\rm qj}$ and $\delta S_{\rm nH}$ in function of the entanglement $S_{e}$ obtained with these trajectories
Appendix C Quantum jumps from highly entangled initial states
In this Appendix we provide evidence to support the conjecture that highly entangled states are more fragile to measurements than weakly entangled ones. To this extent, we consider the monitored SSH chain starting from an initial condition corresponding to the long-time limit of the associated non-Hermitian Hamiltonian, which is known to support volume-law entangled states for small values of $\gamma$. In Fig. 13 (a) we show a sample of trajectories for the entanglement entropy, all converging towards a steady-state value with low entanglement (and equal to the steady-state reached from a product state initial condition). In other words, the system under monitoring cannot sustain volume-law entanglement. In panel (b) we plot the joint distribution of entanglement loss $P(\Delta S_{\rm qj},S)$, which now displays a broadening of its tails indicating that the role of jumps become more relevant. In particular we see that the probability of a large entanglement loss due to atypical jumps increases, which explains why high entanglement cannot be preserved. In panels (c-d) we present the same analysis given in the main text on the first moments of the gain/loss joint distributions, $\overline{\Delta S_{\rm qj}},\overline{\delta S_{\rm nH}}$, but for this particular initial condition. We note that in the region of low entanglement, for $S\leq 2$ (see dashed line) which contains the steady-state (see panel a), the average entanglement gain from non-Hermitian dynamics is essentially independent from subsystem size $\ell$, a behavior similar to what observed in Fig. 9 for a different initial condition. However, as the system dynamics explores larger values of entanglement entropy ($S\geq 2$) throughout its evolution we see important subsystem size effects in $\overline{\delta S_{\rm nH}}$, which would be hidden otherwise starting from a low entangled initial condition.
Appendix D Stochastic model for the SSH setup
In this Appendix we present the results obtained for the monitored SSH model using the classical random-walk with resetting for entanglement dynamics. To obtain these results we proceed as discussed in the main text and summarized in Fig. 1, namely we first compute numerically the joint gain/loss probabilities $P(\delta S_{\rm nH},S),P(\Delta S_{\rm qj},S)$ and then we feed them into the classical master equation to build the associated stochastic process. In Fig. 14 we plot the dynamics of the average entanglement entropy obtained from the random walk (RW) with drift and resetting and compare it with the exact numerical results of the quantum trajectories. As for the monitored Ising chain discussed in Sec. II, the agreement is excellent for both weak and strong monitoring rates $\gamma$, confirming that the classical model is able to capture the entanglement transition also in the monitored SSH chain.
Figure 14: Entanglement dynamics as a classical random walk for the SSH model — Comparison between the dynamics of the average entanglement entropy obtained from the exact quantum jump dynamics (QJ) or the phenomenological classical random walk (RW).
Appendix E Free Fermion Techniques with Quantum Jumps
@@
This Section provides details on solving the stochastic Schrödinger dynamics using free-fermion techniques. Since the Hamiltonian we consider for both models is quadratic, and the monitoring process preserves Gaussianity, the state $|\psi\rangle$ is entirely determined by the 2-point correlation matrix $C$ thanks to Wick’s theorem. We can either store the state in this correlation matrix or use another representation that uses the fact that we consider pure states.
E.1 Correlation matrix
@@
where $C_{m,n}(t)=\langle\Psi(t)|\mathbf{c}_{m}^{\dagger}\mathbf{c}_{n}|\Psi(t)\rangle$ is a matrix of size $2L\times 2L$ and $\mathbf{c}^{T}=(c_{1}\leavevmode\nobreak\ c_{2}\leavevmode\nobreak\ ...\leavevmode\nobreak\ c_{L}\leavevmode\nobreak\ c_{1}^{\dagger}\leavevmode\nobreak\ c_{2}^{\dagger}\leavevmode\nobreak\ ...\leavevmode\nobreak\ c_{L}^{\dagger})$. We get for the deterministic part when measuring the local density, the equation Turkeshi et al. (2022a)
@@
The von Neumann entanglement of a subsystem of size $\ell$ can be extracted through the Majorana fermions correlation matrix Turkeshi et al. (2022a); Vidal et al. (2003). Effectively we get it by diagonalising the following matrix
@@
In the algorithm used in this work, the goal is to find the time at which the jump happens, thus we need to be able to evaluate the decay of the norm during the non-Hermitian evolution to find this time. To do so, we solve numerically the equation
@@
When the model presents a $U(1)$ symmetry it is simpler because the conservation of the number of particles reduces by a factor of 2 the size of the correlation matrix to consider. Indeed, we have to consider only the matrix $G_{m,n}(t)=\langle\psi(t)|\mathbf{c}_{m}^{\dagger}\mathbf{c}_{n}|\psi(t)\rangle$ with $\mathbf{c}^{T}=(c_{1}\leavevmode\nobreak\ c_{2}\leavevmode\nobreak\ ...\leavevmode\nobreak\ c_{L})$. Here we gives the equation that can be used for the SSH model and for convenience we will use a basis respecting the site $A$ and $B$ of the model, thus we consider the basis where $c_{2k-1}=c_{A,k}$ and $c_{2k}=c_{B,k}$ for $21\leq k\leq L/$ 2. In that case we get for the deterministic part (when measuring $n_{A,i}$ and $1-n_{B,i}$), the equation Turkeshi et al. (2022a)

### v3

@@ Appendix E Monitored Ising Chain at Large Field
Figure 12: Steady-State Entanglement Balance for the Monitored Ising chain with large transverse field $h>1$: Average entanglement gain $\overline{\delta S_{\mathrm{nH}}}$ and loss $\overline{\delta S_{\mathrm{qj}}}/\overline{\tau}$, as a function of the average entanglement entropy S and different subsystem sizes $\ell$. Panel (a): the gain contribution due to the non-Hermitian dynamics acquires a non-trivial $\ell$ dependence scaling logarithmically (see inset). On the other hand, the QJ loss term scales also logarithmically which leads to a sub-volume steady-state entanglement absent in the no-click limit. For even larger $h$ (Panel (b), $h=4$), the non-Hermitian gain term becomes $\ell$ independent, as in the no-click limit. At the same time, the QJ retains a non-trivial scaling (inset), leading to the observed logarithmic scaling.
@@
In this Appendix we discuss a particular feature of the monitored Ising chain phase diagram. Indeed, when considering the region of weak monitoring and large transverse field, deviations from the no-click limit were reported Paviglianiti and Silva (2023); Turkeshi et al. (2021). In particular, we fix here $\gamma=0.2$ and scan the transverse field for increasing values of $h>1$, where the no-click evolution predicts an area-law for the entanglement entropy but the QJ dynamics still showed a logarithmic scaling. In Fig. 12 (a-b), we repeat our analysis of average entanglement gain and loss for different subsystem sizes $\ell$. For $h=1.0$, we again observe strong $\ell$ dependence in the QJ contribution and the non-Hermitian one. Our steady-state condition Eq. (21) predicts a logarithmic scaling for the entanglement entropy in agreement with the QJ simulation. Moreover like in the weak transverse field limit by breaking down the entanglement content into gain and loss, we notice that this log-phase is due to a combined effect of logarithmic scale of both : the quantum jumps loss and the renormalized non-Hermitian gain. Upon increasing the transverse field to $h=4$, we see that the entanglement gain due to the non- Hermitian dynamics becomes essentially $\ell$ independent, in agreement with the no-click limit. On the other hand, the loss due to QJs scales logarithmically with $\ell$ (see inset), which results in a steady-state entanglement showing a log-scaling. This logarithmic phase, as for the weak monitoring of the SSH model, is solely due to the QJs.
@@
This analysis therefore highlights two mechanisms at play behind the deviations from the no-click limit observed in the monitored Ising chain phase diagram for $h>1$. At moderate fields there is a non-trivial renormalization of the non-Hermitian dynamics due to QJs, that leads to a log-scaling of its average gain contribution, as in the weak-monitoring regime of the model. At larger fields instead this renormalisation is washed away, the non-Hermitian dynamics behaves as in the no-click limit and the logarithmic scaling of the entropy arises only from the jumps. While our numerical results does not allow to conclude about the large system size limit, the above analysis suggests that, in absence of a non-trivial non-Hermitian dynamics, the large field log-phase might eventually saturate at large system sizes into an area law, similarly to the weak monitoring phase of the SSH model.
Appendix F Free Fermion Techniques with Quantum Jumps
@@
This Section provides details on solving the stochastic Schrödinger dynamics using free-fermion techniques. Since the Hamiltonian we consider for both models is quadratic, and the monitoring process preserves Gaussianity, the state $|\psi\rangle$ is entirely determined by the 2-point correlation matrix $C$ thanks to Wick’s theorem. We can either store the state in this correlation matrix or use another representation based on pure states.
F.1 Correlation matrix
@@
where $C_{m,n}(t)=\langle\Psi(t)|\mathbf{c}_{m}^{\dagger}\mathbf{c}_{n}|\Psi(t)\rangle$ is a matrix of size $2L\times 2L$ and $\mathbf{c}^{T}=(c_{1}\leavevmode\nobreak\ c_{2}\leavevmode\nobreak\ ...\leavevmode\nobreak\ c_{L}\leavevmode\nobreak\ c_{1}^{\dagger}\leavevmode\nobreak\ c_{2}^{\dagger}\leavevmode\nobreak\ ...\leavevmode\nobreak\ c_{L}^{\dagger})$. We get for the deterministic (non-Hermitian) part when measuring the local density, the equation Turkeshi et al. (2022)
@@
The von Neumann entanglement of a subsystem of size $\ell$ can be extracted through the Majorana fermions correlation matrix Turkeshi et al. (2022); Vidal et al. (2003). Effectively we get it by diagonalising the following matrix
@@
In this work we implement the sampling of the waiting-time distribution, therefore the goal is to find the time at which the jump happens. Thus we need to be able to evaluate the decay of the norm during the non-Hermitian evolution and find the time for the next jump. To do so, we solve numerically the equation
@@
A simplification arises when the model presents a $U(1)$ symmetry because the conservation of the number of particles reduces by a factor of 2 the size of the correlation matrix to consider. Indeed, we have to consider only the matrix $G_{m,n}(t)=\langle\psi(t)|\mathbf{c}_{m}^{\dagger}\mathbf{c}_{n}|\psi(t)\rangle$ with $\mathbf{c}^{T}=(c_{1}\leavevmode\nobreak\ c_{2}\leavevmode\nobreak\ ...\leavevmode\nobreak\ c_{L})$. Here we give the equation that can be used for the SSH model and for convenience we will use a basis respecting the site $A$ and $B$ of the model, thus we consider the basis where $c_{2k-1}=c_{A,k}$ and $c_{2k}=c_{B,k}$ for $21\leq k\leq L/$ 2. In that case we get for the deterministic part (when measuring $n_{A,i}$ and $1-n_{B,i}$), the equation Turkeshi et al. (2022)